OpenStudy (fluttershyk):

HELP MEEE PLEASE I WILL MEDAL A researcher tagged 40 giant turtles in 2008. In 2009, she returned and captured 50 turtles, 16 of which were tagged. In 2010, she returned again and captured 30 turtles, 8 of which were tagged. A. Use a proportion to estimate the turtle population for 2009 B. Use a proportion to estimate the turtle population for 2010.

OpenStudy (mysticfury):

is this the whole question?

OpenStudy (bmk614):

First you need to set up a proportion. \[\frac{ 16 }{ 50 }=\frac{ 40 }{ x }\] Then solve for x.

OpenStudy (fluttershyk):

yes that's the whole question @MysticFury

OpenStudy (mysticfury):

oh i was reading it wrong XD

OpenStudy (fluttershyk):

lol

OpenStudy (fluttershyk):

you copy and pasted that @smartnerd1111

OpenStudy (smartnerd1111):

what did i copy and paste

OpenStudy (fluttershyk):

u deleated it

OpenStudy (smartnerd1111):

lol

OpenStudy (fluttershyk):

@rebeccaxhawaii please help I think I know the first one but not the second

OpenStudy (smartnerd1111):

We know that there are 40 turtles tagged right.

OpenStudy (fluttershyk):

yeah the first one is 16/50 = 40/x 16x = 2000 x = 125 but I don't know the second

rebeccaxhawaii (rebeccaxhawaii):

http://openstudy.com/study#/updates/51b3de2ee4b05b167ed3a890

OpenStudy (smartnerd1111):

(P/40 = 30/8) solve p

OpenStudy (smartnerd1111):

rebecca we shouldnt just reference her ot an old openstudy question with the same question asked we should actually help her so next time she would not need to ask a simailr question

OpenStudy (fluttershyk):

ok now im even more confused one somebody just tell me before I bust

OpenStudy (smartnerd1111):

(P/40 = 30/8) solve p what divided by 4 equals 30/8

OpenStudy (smartnerd1111):

40 divided by 8 is 5 so multiply 30 times 5 to get p and that is the proportion of 2010